# Infinite-time Mean Field FBSDEs and Viscosity Solutions to Elliptic Master Equations

### v2

@@ 3 Solutions to infinite-time FBSDEs
$\left.\left.+\hat{Y}_{t}(G(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})-G(t,\omega,X_{t}^{\prime},Y_{t}^{\prime},{\cal L}_{X_{t}^{\prime}}))\right)dt\right]$
$\leq$ $-\kappa\mathbb{E}\left[\int_{0}^{T_{i}}e^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right].$
Letting $T_{i}\to\infty$, we get that
$\|\hat{X}\|_{K}^{2}=\|\hat{Y}\|_{K}^{2}=0,$ (3.8)
and hence we complete the proof of the uniqueness.
Next, we prove the existence of solutions, for this purpose, we use the continuity method. We study the following family of infinite-time FBSDEs parametrized by $\lambda\in[0,1]$,
$\begin{cases}\begin{aligned} \mathrm{d}X_{t}^{\lambda}=&\left[\lambda G(t,\omega,X_{t}^{\lambda},Y_{t}^{\lambda},\mathcal{L}_{X_{t}^{\lambda}})\right.\left.-\kappa(1-\lambda)Y_{t}^{\lambda}+\phi_{t}(\omega)\right]\mathrm{d}t+\sigma\mathrm{d}B_{t},\end{aligned}\\ \begin{aligned} \mathrm{d}Y_{t}^{\lambda}=&-\left[\lambda F(t,\omega,X_{t}^{\lambda},Y_{t}^{\lambda},\mathcal{L}_{X_{t}^{\lambda}})\right.\left.+\kappa(1-\lambda)X_{t}^{\lambda}+\psi_{t}(\omega)\right]\mathrm{d}t+Z_{t}^{\lambda}\mathrm{d}B_{t},\end{aligned}\\ \begin{aligned} &X_{0}^{\lambda}=\xi,\quad(X_{t}^{\lambda},Y_{t}^{\lambda},Z_{t}^{\lambda})\in L_{K}^{2},\end{aligned}\end{cases}$ (3.9)
where $\phi,\psi$ are two arbitrary processes in $L_{K}^{2}$. Note that when $\lambda=1,\phi=\psi=0$, (3.9) becomes (3.1), and when $\lambda=0$, (3.9) becomes
$\begin{cases}&\mathrm{d}X_{t}^{0}=(-\kappa Y_{t}^{0}+\phi_{t}(\omega))\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ &\mathrm{d}Y_{t}^{0}=-(\kappa X_{t}^{0}+\psi_{t}(\omega))\mathrm{d}t+Z_{t}^{0}\mathrm{d}B_{t},\\ &X_{0}^{0}=\xi.\end{cases}$ (3.10)
It has been proved in ([1], Lemma 2.1) that (3.10) has unique solution $(X^{0},Y^{0},Z^{0})\in L_{K}^{2}$.
Now, suppose for some $\lambda_{0}\in[0,1)$, we have that for any ${\cal F}_{0}$-measurable square integrable random variable $\xi$ and $\phi,\psi\in L^{2}_{K}$, (3.9) has a unique solution $(X^{\lambda_{0}},Y^{\lambda_{0}},Z^{\lambda_{0}})$ in $L^{2}_{K}$. We try to find a constant $\delta_{0}$, such that for any $\delta\in[0,\delta_{0}]$, the FBSDE(3.9) has a unique solution $(X^{\lambda_{0}+\delta},Y^{\lambda_{0}+\delta},Z^{\lambda_{0}+\delta})$ in $L^{2}_{K}$ for any given $\xi,\phi,\psi$. To do this, we consider the following FBSDE:
$\begin{cases}\begin{aligned} \mathrm{d}X_{t}={}&\bigl[\lambda_{0}G(t,\omega,X_{t},Y_{t},\mathcal{L}_{X_{t}})-\kappa(1-\lambda_{0})Y_{t}\\ &+\delta\bigl(G(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})+\kappa y_{t}\bigr)+\phi_{t}(\omega)\bigr]\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}={}&-\bigl[\lambda_{0}F(t,\omega,X_{t},Y_{t},\mathcal{L}_{X_{t}})+\kappa(1-\lambda_{0})X_{t}\\ &+\delta\bigl(F(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})-\kappa x_{t}\bigr)+\psi_{t}(\omega)\bigr]\mathrm{d}t+Z_{t}\mathrm{d}B_{t},\\ X_{0}={}&\xi.\end{aligned}\end{cases}$ (3.11)
For any pair $(x_{t},y_{t})$ in $L_{K}^{2}$, we have $G(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})$ and $F(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})$ belong to $L_{K}^{2}$ according to our hypothesis, then the FBSDE (3.11) admits a unique solution $(X,Y,Z)$ in $L_{K}^{2}$. We can define a map $\Phi$ through
$\Phi:(x,y)\mapsto(X,Y).$ (3.12)
We then prove that the map $\Phi$ is a contraction on $L_{K}^{2}$.
Take another pair $(x_{t}^{\prime},y_{t}^{\prime})$ in $L_{K}^{2}$ and its image $(X_{t}^{\prime},Y_{t}^{\prime})$. Denote
$U_{t}=(X_{t},Y_{t},{\cal L}_{X_{t}}),\quad u_{t}=(x_{t},y_{t},{\cal L}_{x_{t}}),$ (3.13)
$\hat{X}_{t}=X_{t}-X_{t}^{\prime},\quad\hat{Y}_{t}=Y_{t}-Y_{t}^{\prime},$
$\hat{x}_{t}=x_{t}-x_{t}^{\prime},\quad\hat{y}_{t}=y_{t}-y_{t}^{\prime}.$
Applying Itô’s formula to $\mathrm{e}^{-Kt}\hat{X}_{t}\hat{Y}_{t}$, we get that
$\left[\mathrm{e}^{-KT}\hat{X}_{T}\hat{Y}_{T}\right]$ (3.14)
$=\lambda_{0}\mathbb{E}\biggl[\int_{0}^{T}\mathrm{e}^{-Kt}\biggl(-K\hat{X}_{t}\hat{Y}_{t}-\hat{X}_{t}\bigl(F(t,\omega,U_{t})-F(t,\omega,U_{t}^{\prime})\bigr)\quad+\hat{Y}_{t}\bigl(G(t,\omega,U_{t})-G(t,\omega,U_{t}^{\prime})\bigr)\biggr)\mathrm{d}t\biggr]$
$-\kappa(1-\lambda_{0})\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right]$
$-(K-\lambda_{0}K)\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\hat{X}_{t}\hat{Y}_{t}\mathrm{d}t\right]$
$+\kappa\delta\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{X}_{t}\hat{x}_{t}+\hat{Y}_{t}\hat{y}_{t}\right)\mathrm{d}t\right]$
$+\delta\mathbb{E}\biggl[\int_{0}^{T}\mathrm{e}^{-Kt}\biggl(-\hat{X}_{t}\bigl(F(t,\omega,u_{t})-F(t,\omega,u_{t}^{\prime})\bigr)\quad+\hat{Y}_{t}\bigl(G(t,\omega,u_{t})-G(t,\omega,u_{t}^{\prime})\bigr)\biggr)\mathrm{d}t\biggr].$
It holds that
$\mathbb{E}\biggl[-K\hat{X}_{t}\hat{Y}_{t}-\hat{X}_{t}\bigl(F(t,\omega,U_{t})-F(t,\omega,U_{t}^{\prime})\bigr)+\hat{Y}_{t}\bigl(G(t,\omega,U_{t})-G(t,\omega,U_{t}^{\prime})\bigr)\biggr]$ (3.15)
$\leq-\kappa\mathbb{E}\left[\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right]$
and
$\mathbb{E}\left[-\hat{X}_{t}\bigl(F(t,\omega,u_{t})-F(t,\omega,u_{t}^{\prime})\bigr)+\hat{Y}_{t}\bigl(G(t,\omega,u_{t})-G(t,\omega,u_{t}^{\prime})\bigr)\right]$ (3.16)
$\leq$ $\mathbb{E}\left[|\hat{X}_{t}|\cdot|F(t,\omega,u_{t})-F(t,\omega,u_{t}^{\prime})|+|\hat{Y}_{t}|\cdot|G(t,\omega,u_{t})-G(t,\omega,u_{t}^{\prime})|\right]$
$\leq$ $\frac{3\ell}{2}\mathbb{E}\left[|\hat{X}_{t}|^{2}+|\hat{Y}_{t}|^{2}\right]+2\ell\mathbb{E}\left[|\hat{x}_{t}|^{2}+|\hat{y}_{t}|^{2}\right].$
So we have
$\mathbb{E}\left[\mathrm{e}^{-KT}\hat{X}_{T}\hat{Y}_{T}\right]\leq$ $-\left(\kappa-\frac{K}{2}-\frac{\kappa\delta+3l\delta}{2}\right)\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right]$ (3.17)
$+\frac{\kappa\delta+4l\delta}{2}\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{x}_{t}^{2}+\hat{y}_{t}^{2}\right)\mathrm{d}t\right].$
We take
$\delta_{0}=\frac{2\kappa-K}{3\kappa+11\ell}$ (3.18)
and choose a sequence of $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-KT_{i}}\hat{X}_{T_{i}}\hat{Y}_{T_{i}}\right]\rightarrow 0.$ (3.19)
For any $\delta\in[0,\delta_{0}]$, we have
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right]\leq\frac{1}{2}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-Kt}\left(\hat{x}_{t}^{2}+\hat{y}_{t}^{2}\right)\mathrm{d}t\right].$ (3.20)
Therefore $\Phi$ is a contraction.
By repeating this procedure for $1/\delta_{0}]$ many times, we conclude that there exists a solution to (3.9) with $\lambda=1$. In particular, we get a $L_{K}^{2}$ solution to (3.1).
@@ 4 Properties of mean field FBSDEs
In this section, we investigate the following mean-field FBSDEs:
@@
To obtain further properties of the above two equations, we require the following assumptions:
@@
where $\hat{X}\triangleq X-X^{\prime},\hat{Y}\triangleq Y-Y^{\prime}$. Since $\kappa-C_{0}>r/2$, the conditions in Assumption 3.1 (iii) is satisfied. Then FBSDE (4.1) has a unique solution in $L_{r}^{2}$.
@@
Now we complete the proof.
Benefiting from the continuous dependence on initial values of the solutions to the equations that we previously proved, we can demonstrate the flow property of the mean field FBSDEs. This precisely reflects the essence of the mean field game: the aggregation of innumerable representative players constitutes the equilibrium state of the system.
@@
and similarly define $Y$ and $Z$ in the same manner. Then $(X,Y,Z)\in L_{r}^{2}$.
@@
Step 2. In the general case, let $\xi_{n}\in\mathbb{L}({\cal F}_{0})$ be a discrete approximation of $\xi$ such that $\mathbb{E}\left[\left\lvert\xi_{n}-\xi\right\rvert^{2}\right]\rightarrow 0$. By Step 1, we have
@@
5 Viscosity solution to master equation
In this section, we will prove the main result of this paper: the value function (1.10) is the viscosity solution to the master equation (1.5).
5.1 Differentiability of mean-field FBSDEs
@@
(ii) There exist constants $\lambda_{1},\lambda_{2}>0$ such that $\lambda_{1}+\lambda_{2}<-r/2$ and
$\partial_{yy}H(x,\mu,y)\leq-\lambda_{1},\quad\partial_{xx}H(x,\mu,y)\geq\lambda_{1},\quad|\partial_{xy}H(x,\mu,y)|\leq\lambda_{2}.$ (5.1)
(iii) There exist a constant $\lambda_{3}>0$ such that
$|\partial_{xx}H(x,\mu,y)|\leq\lambda_{3},\quad|\partial_{yy}H(x,\mu,y)|\leq\lambda_{3}.$ (5.2)
@@
Under Assumption 4.1 and 5.1, it’s clear that the above FBSDE satisfies all conditions in Assumption 3.1, so it admits a unique solution in $L_{r}^{2}$. The following theorem tells us that $(\nabla X^{x,\xi},\nabla Y^{x,\xi})$ can be viewed as the derivative of $(X^{x,\xi},Y^{x,\xi})$ with respect to $x$.
@@
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(X^{x+\delta,\xi}-X^{x,\xi}\right)-\nabla X^{x,\xi}\right\rVert_{r}\to 0,$ (5.4)
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(Y^{x+\delta,\xi}-Y^{x,\xi}\right)-\nabla Y^{x,\xi}\right\rVert_{r}\to 0.$
@@
$\nabla X^{\delta}\triangleq\frac{\Delta X^{\delta}}{\delta},\quad\nabla Y^{\delta}\triangleq\frac{\Delta Y^{\delta}}{\delta},\quad\nabla Z^{\delta}\triangleq\frac{\Delta Z^{\delta}}{\delta}.$ (5.7)
@@
Then we have $(X^{\delta}_{t},Y^{\delta}_{t},Z^{\delta}_{t})$ are uniformly bounded in $L_{r}^{2}$, and they satisfy the following FBSDE:
@@
From the a priori estimate, we know that $Y_{t}^{\delta}$ is uniformly bounded in $L_{r}^{2}$, and $|A_{t}^{\delta}|\leq 2\lambda_{3}$, then there exists a constant $C>0$, such that
@@
as $\delta\to 0$. This shows that under measure $\mathbb{R}$, $|A^{\delta}|^{2}$ converges in measure to 0. Since $(\nabla X^{x,\xi})^{2}$ is integrable under $\mathbb{Q}$, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ converges in measure to 0. Finally, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ is dominated by an integrable function $4\lambda_{3}^{2}(\nabla X^{x,\xi})^{2}$ under measure Q, and by the Dominated Convergence Theorem, we have
@@
Now we finish the proof.
Now we suppose $F(x,\mu,y):\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}\rightarrow\mathbb{R}$ satisfies the following conditions:
@@
where $\xi\in\mathbb{L}^{2}({\cal F}_{0})$ and ${\cal L}_{\xi}=\mu$. Since we have already established the uniqueness in distribution for the FBSDEs (see [14]), $U$ is well-defined.
@@
5.2 Main results
@@
Under Assumption 4.1 and 5.1, the value function satisfies
@@
It’s easy to verify that $f(x,\mu,\hat{\alpha}(x,y))$ satisfies all conditions in Lemma 5.3, so $\partial_{x}V(x,\mu)$ exists and
@@
Now we get the desired result.
In preparation for the definition of a viscosity solution of the master equation, we first define the class of test functions used for that purpose.
@@
Now, we present the definition of the viscosity solution for master equation (1.5).
@@
(respectively
$\begin{split}rU(x^{0},\mu^{0})\geq&H(x^{0},\mu^{0},\partial_{x}U(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\cdot\partial_{y}H(\tilde{\xi}^{0},\mu^{0},\partial_{x}U(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (5.44)
@@
Under Assumption 4.1 and 5.1, the value function (1.10) is the viscosity solution to the master equation (1.5).
@@
By the flow property and dynamic programming principle, we have that
@@
Then by the relationship (5.45), we have that

### v3

@@ 2 Preliminaries
We have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big[\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\partial_{xx}U(\sigma_{t}^{1})^{2}\Big](t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big(\tilde{\mathbb{E}}_{{\cal F}_{t}}\big[\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{\sigma}_{t}^{2})^{2}\big]\Big)\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.9)
Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ and the process $(\tilde{X}^{2}_{t},\tilde{b}^{2}_{t},\tilde{\sigma}_{t}^{2})_{0\leq t\leq T}$ is a copy of the process $(X^{2}_{t},b^{2}_{t},\sigma_{t}^{2})_{0\leq t\leq T}$, on a copy $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ of the probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$.
@@ 3 Properties of mean field FBSDEs
In this section, for some $\xi\in\mathbb{L}^{2}({\cal F}_{0})$, we investigate the following mean field FBSDEs:
@@
$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x,\xi}=x.\end{cases}$ (3.2)
In [14], we provided sufficient conditions on $b(x,\mu,a)$ and $f(x,\mu,a)$ to ensure the above two FBSDEs admit unique solutions. To obtain further properties of these solutions , we require the following assumptions on $H$:
@@
where $\hat{X}\triangleq X-X^{\prime},\hat{Y}\triangleq Y-Y^{\prime}$. Since $\kappa-C_{0}>r/2$, the conditions in Assumption 2.1 (iii) is satisfied. Then, FBSDE (3.1) has a unique solution in $L_{r}^{2}$.
@@
Now, we complete the proof.
Benefiting from the continuous dependence on initial values of the solutions to the mean field FBSDEs, we can demonstrate the flow property. This precisely reflects the essence of the mean field game: the aggregation of innumerable representative players constitutes the equilibrium state of the system.
@@
and similarly define $Y$ and $Z$ in the same manner. Then, we know $(X,Y,Z)\in L_{r}^{2}$.
@@
Step 2. In the general case, let $\xi_{n}\in\mathbb{L}^{2}({\cal F}_{0})$ be a discrete approximation of $\xi$ such that $\mathbb{E}\left[\left\lvert\xi_{n}-\xi\right\rvert^{2}\right]\rightarrow 0$. By Step 1, we have
@@
4 Solutions to elliptic master equations
In this section, we will first prove the main result of this paper: the value function (1.14) is a viscosity solution to the elliptic master equation (1.9). And then we show the uniqueness of classical solutions to the elliptic master equation under suitable growth conditions and the displacement monotonicity condition.
4.1 Differentiability of mean field FBSDEs
@@
(ii) There exist a constant $\lambda_{1}>r/2$ such that
$\partial_{yy}H(x,\mu,y)\leq-\lambda_{1},\quad\partial_{xx}H(x,\mu,y)\geq\lambda_{1}.$ (4.1)
(iii) There exist a constant $\lambda_{2}>0$ such that
$|\partial_{xx}H(x,\mu,y)|\leq\lambda_{2},\quad|\partial_{yy}H(x,\mu,y)|\leq\lambda_{2},\quad|\partial_{xy}H(x,\mu,y)|\leq\lambda_{2}.$ (4.2)
@@
Under Assumption 3.1 and 4.1, it’s clear that the above FBSDE satisfies all conditions in Assumption 2.1, so it admits a unique solution in $L_{r}^{2}$. The following theorem tells us that $(\nabla X^{x,\xi},\nabla Y^{x,\xi})$ can be viewed as the derivative of $(X^{x,\xi},Y^{x,\xi})$ with respect to $x$.
@@
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(X^{x+\delta,\xi}-X^{x,\xi}\right)-\nabla X^{x,\xi}\right\rVert_{r}=0,$ (4.4)
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(Y^{x+\delta,\xi}-Y^{x,\xi}\right)-\nabla Y^{x,\xi}\right\rVert_{r}=0.$
@@
$\nabla X^{\delta}\triangleq\frac{\Delta X^{\delta}}{\delta},\quad\nabla Y^{\delta}\triangleq\frac{\Delta Y^{\delta}}{\delta},\quad\nabla Z^{\delta}\triangleq\frac{\Delta Z^{\delta}}{\delta},$ (4.7)
@@
Then, we have $(X^{\delta}_{t},Y^{\delta}_{t},Z^{\delta}_{t})$ are uniformly bounded in $L_{r}^{2}$, and they satisfy the following FBSDE:
@@
From the a priori estimate, we know that $Y_{t}^{\delta}$ is uniformly bounded in $L_{r}^{2}$, and $|A_{t}^{\delta}|\leq 2\lambda_{2}$, then there exists a constant $C>0$, such that
@@
as $\delta\to 0$. This shows that under measure $\mathbb{R}$, $|A^{\delta}|^{2}$ converges in measure to 0. Since $(\nabla X^{x,\xi})^{2}$ is integrable under $\mathbb{Q}$, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ converges in measure to 0. Finally, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ is dominated by an integrable function $4\lambda_{2}^{2}(\nabla X^{x,\xi})^{2}$ under measure Q, and by the Dominated Convergence Theorem, we have
@@
Now, we finish the proof.
Now, we consider a function $F(x,\mu,y):\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}\rightarrow\mathbb{R}$ satisfying the following conditions:
@@
where $\xi\in\mathbb{L}^{2}({\cal F}_{0};\mu)$. Since we have already established the uniqueness in distribution for the FBSDEs (see [14]), $U$ is well-defined.
@@
4.2 Viscosity solution to elliptic master equation
@@
Under Assumption 3.1 and 4.1, the value function is continuous and satisfies
@@
It’s easy to verify that $H(x,\mu,y)$ satisfies all conditions in Lemma 4.3, whereas $\partial_{y}H(x,\mu,y)\cdot y$ fails to do so. We therefore require a more delicate analysis. Fix $x\in\mathbb{R},\xi\in\mathbb{L}^{2}({\cal F}_{0})$, we continue to use the notation from the proof of Theorem 4.2. For any $|\delta|<1$, set
$U^{\delta}\triangleq$ $\frac{1}{\delta}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\partial_{y}H(X_{t}^{x+\delta,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x+\delta,\xi})\cdot Y^{x+\delta,\xi}_{t}-\partial_{y}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y^{x,\xi}_{t}\right)\mathrm{d}t\right]$ (4.37)
$=$ $\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y^{\delta}_{t}\partial_{y}H(X_{t}^{x+\delta,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x+\delta,\xi})\mathrm{d}t\right]$
$+\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}Y_{t}^{x,\xi}\left(H_{xy}^{\delta}\cdot\nabla X_{t}^{\delta}+H_{yy}^{\delta}\cdot Y_{t}^{\delta}\right)\mathrm{d}t\right]$
and
$U\triangleq$ $\mathbb{E}\bigg[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y_{t}^{x,\xi}\partial_{y}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})$ (4.38)
$+Y_{t}^{x,\xi}\left(\partial_{xy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla X_{t}^{x,\xi}+\partial_{yy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla Y_{t}^{x,\xi}\right)\mathrm{d}t\bigg].$
On the one hand, by an application of Hölder’s inequality, we can obtain that
$\lim_{\delta\to 0}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y^{\delta}_{t}\partial_{y}H(X_{t}^{x+\delta,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x+\delta,\xi})\mathrm{d}t\right]$ (4.39)
$=$ $\mathbb{E}\bigg[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y_{t}^{x,\xi}\partial_{y}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t\bigg].$
On the other hand, employing a method similar to that used in the proof of Theorem 4.2, which utilizes the Dominated Convergence Theorem, we can obtain
$\lim_{\delta\to 0}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}Y_{t}^{x,\xi}\left(H_{xy}^{\delta}\cdot\nabla X_{t}^{\delta}+H_{yy}^{\delta}\cdot Y_{t}^{\delta}\right)\mathrm{d}t\right]$ (4.40)
$=$ $\mathbb{E}\bigg[\int_{0}^{\infty}\mathrm{e}^{-rt}Y_{t}^{x,\xi}\left(\partial_{xy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla X_{t}^{x,\xi}+\partial_{yy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla Y_{t}^{x,\xi}\right)\mathrm{d}t\bigg].$
Therefore
$\lim_{\delta\to 0}U^{\delta}=U.$ (4.41)
Noting that all the first-order partial derivatives of $f(x,\mu,\hat{\alpha}(x,y))$ have at most linear growth, we obtain from Lemma 4.3 that the value function $V(x,\mu)$ is continuous and $\partial_{x}V(x,\mu)$ exists, which satisfies
@@
Now, we get the desired result.
In preparation for the definition of a viscosity solution of the elliptic master equation, we first define the class of test functions used for that purpose.
@@
Now, we present the definition of the viscosity solution for elliptic master equation (1.9).
@@
(respectively,
$\begin{split}rU(x^{0},\mu^{0})\geq&H(x^{0},\mu^{0},\partial_{x}U(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\cdot\partial_{y}H(\tilde{\xi}^{0},\mu^{0},\partial_{x}U(\tilde{\xi}^{0},\mu^{0}))\right],\end{split}$ (4.49)
@@
Under Assumption 3.1 and 4.1, the value function (1.14) is the viscosity solution to the elliptic master equation (1.9).
@@
By the flow property, we have that
@@
Then, by the relationship (4.50), we have that